# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 12.2 Decision Algorithm and its verification
    GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle)$.
@@
    For each $a\in C\langle E\cup F,\emptyset\rangle$, GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a$.
@@
    For each $a\in C\langle E\cup F,\emptyset\rangle$,
    • $a.equal_{E}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{E}=\mathrm{true}$, or
    • $a.equal_{F}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
@@
Recall that we have run Algorithm CAD on $\Sigma$, $E$, and $F$. By Proposition 4.47, we have obtained in $O(n\,\mathrm{log}\,n)$ time the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
@@
We now continue Examples 3.17, 5.50, 7.71, 10.95. and 12.109. Recall that we run Algorithm NPDFS on the auxiliary dpwpa $AUX[E;F]$, and Algorithm NPDFS outputs true to the question whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Hence Condition (iii) in Lemma 12.112 holds. Consequently, by Lemma 12.112, $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
@@
We now continue Examples 3.18, 5.51, 7.72, 10.96, and 12.110. Recall that we run Algorithm NPDFS on the auxiliary dpwpa $AUX[E;F]$, and Algorithm NPDFS outputs false to the question whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. For this reason Condition (iii) in Lemma 12.112 does not hold. Consequently, by Lemma 12.112, $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
@@ 13 Main Case 2
In this main case, $E$ and $F$ are GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total.
@@
Proof.
Lemma 9.87 and Proposition 9.84, and Proposition 9.82 imply the Theorem. ∎
We present a faster and simpler algorithm for solving our decidability problem. We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.47. We showed that $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class. By direct inspection of $INC(E\cup F,E)$ stored in a red-black tree, and $INC(E\cup F,F)$ stored in a red-black tree we decide in $O(n)$ time whether each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class. Thence we decide in $O(n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Example 3.19 and its sequels, Examples 5.52, 13.117, in addition Example 3.20 and its sequels, Examples 5.53, 13.118. exemplify this case.
@@
We now continue Examples 3.19 and 5.52. Recall that
@@
  • #$\#/_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • \$$\$/_{\Theta\langle E,F\rangle}=\$/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • £$\mathsterling/_{\Theta\langle E,F\rangle}=\mathsterling/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • #$\#/_{\Theta\langle F,E\rangle}=\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • \$$\$/_{\Theta\langle F,E\rangle}\subset\$/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • £$\mathsterling/_{\Theta\langle F,E\rangle}=\mathsterling/_{\Theta\langle{E\cup F},\emptyset\rangle}$, and
@@
We now continue Examples 3.20 and 5.53. Recall that
@@
  • #$\#/_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and #$\#/_{\Theta\langle{F},E\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$, and
@@
Justified by the above observations and by Statement 6.60, we compute as follows. Recall that $NUM(E\cup F,E)$ stored in a red-black tree $RBT3$, and $NUM(E\cup F,F)$ stored in a red-black tree $RBT4$. Alternately in a coordinated way, we traverse the red-black trees $RBT3$ and $RBT4$ in in-order traversal, For each $a\in C\langle E\cup F,\emptyset\rangle$, we read the pair $(a,k)$, where $k=|\{b\in C\langle E,F\rangle\mid b\subseteq a\,\}|$, and the pair $(a,\ell)$, where =|$\ell=|\{b\in C\langle F,E\rangle\mid b\subseteq a\,\}|$. In this way, we decide in $O(n)$ time whether each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class, see Definition 4.26. If the answer is ’yes’, then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, otherwise, $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. ∎
@@ 14 Main Case 3
In this main case, $E$ and $F$ are GTESs over a signature $\Sigma$ such that $\Sigma$ has a symbol of arity at least $2$, and $R\langle E,F\rangle$ or $R\langle F,E\rangle$ are total.
@@
We present a faster and simpler decision algorithm. Without loss of generality, we may assume that $R\langle E,F\rangle$ is total from now on throughout this section. Example 3.21 and its sequels, Examples 5.54, 7.73, 10.100, 14.125, and 14.132 plus Example 3.22, and its sequels, Examples 5.55, 7.74, 10.101, 14.126, and 14.133 present this case. We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.47.
@@ 14.1 Partial Depth First Search on the Auxiliary dpwpa
var $a:C\langle E\cup F,\emptyset\rangle$;
@@
for each vertex $a\in C\langle E\cup F,\emptyset\rangle$ do
@@
{  if $a\not\in C\langle E,F\rangle$ and $a\not\in C\langle F,E\rangle$ then GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. }
$a.visited:=\mathrm{false}$
@@
for each vertex $a\in C\langle E\cup F,\emptyset\rangle$ do
if not $a.equal_{E}$ and not $a.visited$
{ if $a\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$ and $a$ has not been visited yet }
${\bf then}$ ${\bf for}$ ${\bf each}$ vertex $b\in Adj(a)$ do
${\bf if}$ not $b.visited$
${\bf then}$ if not $\mathrm{PDFS-VISIT}(AUX[E;F],b)$ then
@@
${\bf function}\;\mathrm{PDFS-VISIT}(AUX[E;F],a):\;\mathrm{Boolean}$
$a.visited:=\mathrm{true}$;
if $a.keeps_{F}=\mathrm{true}$
@@
for each vertex $b\in Adj(a)$ do
${\bf if}$ not $b.visited$
${\bf then}$ ${\bf if}$ ${\bf not}$ $\mathrm{PDFS-VISIT}(AUX[E;F],b)$
@@
When Algorithm PDFS starts the search from a vertex $a\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$, it does not set the attribute $a.visited$ to true. Consequently if for this vertex $a\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$, there exists a cycle leading from the vertex $a$ to itself, then Algorithm PDFS visits vertex $a$ twice.
@@
[8] Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and let $E$ and $F$ be GTESs over $\Sigma$ such that $R\langle E,F\rangle$ is total. Let $a\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$.
(i) During the run of Algorithm PDFS on $AUX[E;F]$, Function $\mathrm{PDFS-VISIT}$ on $(AUX[E;F],a)$ visits at most once those vertices $b\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from $a$ to $b$.
(ii) During the run of Algorithm PDFS on $AUX[E;F]$, Function $\mathrm{PDFS-VISIT}$ visits at most once those vertices $b\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex $a\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to $b$.
@@
The running time of PDFS is $O(|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|)$ [8]. Recall that $n=size(E)+size(F)$. By Definition 4.26 and Statement 4.45, $|C\langle E\cup F,\emptyset\rangle|\leq|W|\leq n$. By Statement 6.63 $|A[E;F]|\leq n$. Hence $|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|\leq 2n$. Then we have the proposition. ∎
@@
We now continue Examples 3.21, 5.54, 7.73, and 10.100. First Algorithm PDFS decides whether for each vertex $a\in C\langle E\cup F,\emptyset\rangle$, it holds that $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$, that is, $\left(a.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$. Recall that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$. As a result
(#$\left(\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}\;\,\mbox{and}\;\,\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}\right)=\mathrm{true}$.
@@
We now continue Examples 3.22, 5.55, 7.74, and 10.101. Recall that there exist one path of positive length and two cycles in $AUX[E;F]$.
@@
The first cycle: Its length is $1$, and it starts from the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
@@
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
@@
The path of positive length: Its length is $1$, and it starts from the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
@@
\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
@@
The second cycle: Its length is $1$, and it starts from the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
@@
\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
@@
First Algorithm PDFS decides whether for each $a\in C\langle E\cup F,\emptyset\rangle$, it holds that $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$, that is, $\left(a.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$. Observe that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$. Thus
@@
(#$\left(\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}\;\mbox{or}\;\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}\right)=\mathrm{true}$.

### v2

@@ 12.2 Decision Algorithm and its verification
    For each ${b}\in C\langle E\cup F,\emptyset\rangle$, GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant b{b}.
@@
    For each ${b}\in C\langle E\cup F,\emptyset\rangle$,
    • (a)
      ${b}.equal_{E}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{E}=\mathrm{true}$, or
    • (b)
      ${b}.equal_{F}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{F}=\mathrm{true}$.
@@
Recall that we have run Algorithm CAD on $\Sigma$, $E$, and $F$. By Proposition 4.48, we have obtained in $O(n\,\mathrm{log}\,n)$ time the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes ${b}.equal_{E}$, ${b}.equal_{F}$, ${b}.keeps_{E}$, and ${b}.keeps_{F}$.
@@
We now continue Examples 3.18, 5.51, 7.72, 10.105. and 12.119. Recall that we run Algorithm NPDFS on the auxiliary dpwpa $AUX[E;F]$, and Algorithm NPDFS outputs true to the question whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Hence Condition (iii) in Lemma 12.122 holds. Consequently, by Lemma 12.122, $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
@@
We now continue Examples 3.19, 5.52, 7.73, 10.106, and 12.120. Recall that we run Algorithm NPDFS on the auxiliary dpwpa $AUX[E;F]$, and Algorithm NPDFS outputs false to the question whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. For this reason Condition (iii) in Lemma 12.122 does not hold. Consequently, by Lemma 12.122, $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
@@ 13 Main Case 2
In this main case, $E$ and $F$ are GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total. We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.48. We showed that $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class. By direct inspection of $INC(E\cup F,E)$ stored in a red-black tree and $INC(E\cup F,F)$ stored in a red-black tree we decide in $O(n)$ time whether each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class. Thence we decide in $O(n\,\mathrm{log}\,n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Example 3.20 and its sequels, Examples 5.53 and 13.127, in addition Example 3.21 and its sequels, Examples 5.54 and 13.128 exemplify this case. First we note that we can also apply the decision algorithm of Champavère et al [10]. Lemma 9.97 and Proposition 9.94, and Proposition 9.92 imply the following theorem.
@@
We present a faster and simpler algorithm for solving our decidability problem.
@@
We now continue Examples 3.20 and 5.53. Recall that
  • 1.
@@
  • 2.
    $_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • 3.
    $\mathdollar/_{\Theta\langle E,F\rangle}=\mathdollar/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • 4.
    $\mathsterling/_{\Theta\langle E,F\rangle}=\mathsterling/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • 5.
    $_{\Theta\langle F,E\rangle}=\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • 6.
    $\mathdollar/_{\Theta\langle F,E\rangle}\subset\mathdollar/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • 7.
    $\mathsterling/_{\Theta\langle F,E\rangle}=\mathsterling/_{\Theta\langle{E\cup F},\emptyset\rangle}$, and
  • 8.
@@
We now continue Examples 3.21 and 5.54. Recall that
  • 1.
@@
  • 3.
    $_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and $_{\Theta\langle{F},E\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$, and
  • 4.
@@
Justified by the above observations and by Statement 6.61, we compute as follows. Recall that $NUM(E\cup F,E)$ stored in a red-black tree $RBT3$, and $NUM(E\cup F,F)$ stored in a red-black tree $RBT4$. Alternately in a coordinated way, we traverse the red-black trees $RBT3$ and $RBT4$ in in-order traversal, For each ${b}\in C\langle E\cup F,\emptyset\rangle$, we read the pair (b,k)({b},k), where $k=|\{{c}\in C\langle E,F\rangle\mid{c}\subseteq{b}\,\}|$, and the pair $({b},\ell)$, where $\ell=|\{{c}\in C\langle F,E\rangle\mid{c}\subseteq{b}\,\}|$. In this way, we decide in $O(n)$ time whether each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class, see Definition 4.27. If the answer is “yes”, then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, otherwise, $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. ∎
@@ 14 Main Case 3
In this main case, $E$ and $F$ are GTESs over a signature $\Sigma$ such that $\Sigma$ has a symbol of arity at least $2$, and $R\langle E,F\rangle$ or $R\langle F,E\rangle$ are total. Without loss of generality, we may assume that $R\langle E,F\rangle$ is total from now on throughout this section. Example 3.22 and its sequels, Examples 5.55, 7.74, 10.110, 14.135, and 14.142 plus Example 3.23, and its sequels, Examples 5.56, 7.75, 10.111, 14.136, and 14.143 present this case. We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.48. First we note that we can also apply the decision algorithm of Champavère et al [10].
@@
We present a faster and simpler decision algorithm.
@@ 14.1 Partial Depth First Search on the Auxiliary dpwpa
var ${b}:C\langle E\cup F,\emptyset\rangle$;
@@
for each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ do
@@
{  if ${b}\not\in C\langle E,F\rangle$ and ${b}\not\in C\langle F,E\rangle$ then GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. }
${b}.visited:=\mathrm{false}$
@@
for each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ do
if not ${b}.equal_{E}$ and not b.visited{b}.visited
{ if ${b}\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$ and b{b} has not been visited yet }
$\bf then$} $\bf for$} $\bf each$} vertex ${c}\in Adj({b})$ do
$\bf if$} not c.visited{c}.visited
$\bf then$} if not $\mathrm{PDFS-VISIT}(AUX[E;F],{c})$ then
@@
${\bf function}\;\mathrm{PDFS-VISIT}(AUX[E;F],{b}):\;\mathrm{Boolean}$
${b}.visited:=\mathrm{true}$;
if ${b}.keeps_{F}=\mathrm{true}$
@@
for each vertex ${c}\in Adj({b})$ do
$\bf if$} not c.visited{c}.visited
$\bf then$} $\bf if$} $\bf not$} $\mathrm{PDFS-VISIT}(AUX[E;F],{c})$
@@
When Algorithm PDFS starts the search from a vertex ${b}\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$, it does not set the attribute b.visited{b}.visited to true. Consequently if for this vertex ${b}\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$, there exists a cycle leading from the vertex b{b} to itself, then Algorithm PDFS visits vertex b{b} twice.
@@
[13] Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and let $E$ and $F$ be GTESs over $\Sigma$ such that $R\langle E,F\rangle$ is total. Let ${b}\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$.
(i) During the run of Algorithm PDFS on $AUX[E;F]$, Function $\mathrm{PDFS-VISIT}$ on (AUX[E;F],b)(AUX[E;F],{b}) visits at most once those vertices ${c}\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from b{b} to c{c}.
(ii) During the run of Algorithm PDFS on $AUX[E;F]$, Function $\mathrm{PDFS-VISIT}$ visits at most once those vertices ${c}\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex ${b}\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to c{c}.
@@
The running time of PDFS is $O(|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|)$ [13]. Recall that $n=\mathrm{size}(E)+\mathrm{size}(F)$. By Definition 4.27 and Statement 4.46, $|C\langle E\cup F,\emptyset\rangle|\leq|W|\leq n$. By Statement 6.64 $|A[E;F]|\leq 2n$. Hence $|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|\leq 3n$. Then we have the proposition. ∎
@@
We now continue Examples 3.22, 5.55, 7.74, and 10.110. First Algorithm PDFS decides whether for each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, it holds that ${b}\in C\langle E,F\rangle$ or ${b}\in C\langle F,E\rangle$, that is, $\left({b}.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$. Recall that $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$. As a result
$\left(\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}\;\,\mbox{and}\;\,\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}\right)=\mathrm{true}$.
@@
We now continue Examples 3.23, 5.56, 7.75, and 10.111. Recall that there exist one path of positive length and two cycles in $AUX[E;F]$.
@@
The first cycle: Its length is $1$, and it starts from the vertex $_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to $_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant $_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
@@
$_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$, and $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
@@
The path of positive length: Its length is $1$, and it starts from the vertex $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to $_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant $_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
@@
$\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
@@
The second cycle: Its length is $1$, and it starts from the vertex $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
@@
$\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
@@
First Algorithm PDFS decides whether for each ${b}\in C\langle E\cup F,\emptyset\rangle$, it holds that ${b}\in C\langle E,F\rangle$ or ${b}\in C\langle F,E\rangle$, that is, $\left({b}.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$. Observe that $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$. Thus
@@
$\left(\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}\;\mbox{or}\;\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}\right)=\mathrm{true}$.